OpenStudy (anonymous):

Solve for x. x/4 +8 =5 x = 52 x = -12 x = -15 x = -9

OpenStudy (anonymous):

will medal if correct

OpenStudy (anonymous):

x = -12

OpenStudy (anonymous):

can you help me with more question

OpenStudy (anonymous):

-12/4 = -3 -3+8 = 5

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

1. Addison worked 10 more hours this week than last week. In total, she worked 80 hours this week and last week. Using the equation 2h + 10 = 80, how many hours (h) did Addison work last week? 35 8 68 70

OpenStudy (anonymous):

70

OpenStudy (anonymous):

not correct uwu

OpenStudy (anonymous):

Which value for x makes the sentence true? 10x = 3x + 14 x = 21 x = 1 x = 2 x = 7

OpenStudy (anonymous):

x = 2 10x2 = 20 3x2 = 6, 6+14= 20

OpenStudy (anonymous):

Sorry for the last the one

OpenStudy (anonymous):

it's pl sp it's 21 this one Which value for x makes the sentence true? 10x = 3x + 14 x = 21 x = 1 x = 2 x = 7

OpenStudy (anonymous):

x = 2

OpenStudy (anonymous):

come back XD

OpenStudy (anonymous):

Which value of x makes the following equation true? 3(x - 5) - 12 = 2(x - 6) + 4 25 13 19 14 im still trying to finish

OpenStudy (anonymous):

x=19

OpenStudy (anonymous):

ok Which of the following best describes the solution to the equation below? no real solutions exactly one real solution, infinite real solutions exactly one real solution,

OpenStudy (anonymous):

Wheres the equation?

OpenStudy (anonymous):

30+6x =9x

OpenStudy (anonymous):

Sorry, I'm not quite sure. I have a vague idea of what it might be, but you don't want to take that chance.

OpenStudy (anonymous):

i got a better chance with yours

OpenStudy (anonymous):

Which value of x makes the following equation true? 3(x - 2) + 36 = -3x

OpenStudy (anonymous):

-5 -4 -4 1/3 5

OpenStudy (anonymous):

x= -5

OpenStudy (anonymous):

correct and thank you Which value of x makes the following equation true? 2(x + 8) = x + 19 -5 11 3 19

OpenStudy (anonymous):

x= 3

OpenStudy (anonymous):

correct again XD . Which value of x makes the following equation true? 7.7(x - 4) = -0.3x + 33.2 11 10 7 8

OpenStudy (anonymous):

x= 8

OpenStudy (anonymous):

correct Which value for x makes the sentence true? 15x - 3 = 7x + 13 2 5 -1 4

OpenStudy (anonymous):

x= 2

OpenStudy (anonymous):

Which value for x makes the sentence true? 11x - 4 = 4x + 10 4 2 -2 6

OpenStudy (anonymous):

x= 2

OpenStudy (anonymous):

thank you once again i won't get chewed out XD Addison worked 8 more hours this week than last week. In total, she worked 66 hours this week and last week. Using the equation 2h + 8 = 66, how many hours (h) did Addison work last week? 56 58 29 8

OpenStudy (anonymous):

My guess would be 56, but I'm not sure that's correct

OpenStudy (anonymous):

Susannah has a high-yield savings account into which she deposits $200 each month. At the beginning of the year, Susannah's account balance is $700. In the equation below, x represents the amount of time, and y represents Susannah's account balance. y = $200x + $700 Not including interest, what will her account balance be at the end of the year? $2,700 $3,200 $3,100 $2,400

OpenStudy (anonymous):

$3,100

OpenStudy (anonymous):

Hans is using frequent flier miles to fly to a location 700 miles away, but the airline he is using still charges $15.97 for fees and taxes and $0.14 to redeem each mile. In the equation below, x represents the distance Hans is flying, and y represents the cost of the trip. y = $0.14x + $15.97 How much did Hans pay for his trip? $113.97 $82.03 $106.97 $129.94

OpenStudy (anonymous):

$113.97

OpenStudy (anonymous):

almost done Starla bought cloth napkins at an outlet store for $2.50 each, and she also used a $20 off coupon. In the equation below, x represents the number of napkins, and y represents the cost before tax. y = $2.50x - $20 If Starla paid $35.00 before tax, how many napkins did she buy? 24 14 7 22

OpenStudy (anonymous):

22

OpenStudy (anonymous):

Which value of x makes the following equation true? 3(x - 1) = -2x + 27 6 7 9 10

OpenStudy (anonymous):

crap

OpenStudy (anonymous):

What do you think it is?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

That's not correct, would you like me to show you how to do it or would you like to try it one more time?

OpenStudy (anonymous):

show me pls

OpenStudy (anonymous):

6-1 = 5 5x3 = 15 Now on the other side of the equation -2 x 6 = -12 -12 + 27 = 15 So x = 6 Do you understand?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

I replaced x with 6 in the equation. We wanted both sides of the equation to match. So when both sides were 15, we knew that x = 6

OpenStudy (anonymous):

i understand better

OpenStudy (anonymous):

Anymore questions?

OpenStudy (anonymous):

thank you again and no i will close this question now

OpenStudy (anonymous):

Your welcome, if you need anymore help on something, let me know

OpenStudy (anonymous):

ok and i will XD bya